OpenStudy (anonymous):

3x + 6 + 1 = 8 Would that part of an equation be considered combining like terms?

OpenStudy (saifoo.khan):

3x + 7 = 8 3x = 8-7 3x = 1 x = 1/3

OpenStudy (anonymous):

No, I know the answer and how to solve it, my question was: is the part of the equation i posted considered combining like terms? Then the second part: 3x + 7 = 8 be Distributive Property.